HI i need help with these questions.(Completing the square.)
1.)75x^2+35x-10=0
2.)4x^2-72x+324=0

Found 2 solutions by lynnlo, stanbon:
Answer by lynnlo(4176)

(1)x=1/5,-2/3
(2)x=9

Answer by stanbon(75887)

Completing the square.)
1.)75x^2+35x-10=0
25x^2 + 7x - 5 = 0
----
25x^2 + 7x = 5
25(x^2 + (7/25)x + ?) = 5 + 25*?
--------------
Complete the square:
25(x^2 + 7/25 x + (7/50)^2) = 5 + 25(7/50)^2
--------------
25(x+(7/50))^2 = 5 + 25(49/2500)
---
25(x+(7/50))^2 = 5 + (49/100)
25(x+(7/50))^2 = (5*49/100)
---
(x+(7/50))^2 = (49/500)
------
x+(7/50) = 7/(10sqrt(5))
etc.
===============

2.)4x^2-72x+324=0
4x^2 -72x = -324
----
4(x^2-18x + ?) = -324 + 4*?
Complete the square:
4(x^2-18x+81) = -324 + 4*81
----
4(x-9)^2 = 0
------
x = 9 with multiplicity 2.
Cheers,
Stan H.
===============